OpenStudy (anonymous):

Simplify cot θ tanθ. A.sec θ B.cos θ C.sin θ D.1

OpenStudy (anonymous):

$$\cot \theta=\frac1{\tan \theta}$$Notice they're multiplicative inverses... if you take their product it equals the multiplicative identity, \(1\), just like \(2\times\frac12\)